OpenStudy (anonymous):

An 0.967 mol sample of xenon gas at a temperature of 15.0 °C is found to occupy a volume of 23.7 liters. The pressure of this gas sample is mm Hg.

OpenStudy (somy):

n= 0.967mol T= 15 Celsius = needs to be converted to Kelvin= 15+273= 288 K Volume= 23.7 Liters = needs to be converted to m^3= 23.7/1000= 0.0237m^3 Pressure= ? Formula used here is PV=nRT (R is a constant= 8.31) so we need pressure P= nRT/V P= (0.967*8.31*288) / 0.0237 = 97 649.86329 Pa now we need to convert Pa to mmHg (check out conversion in google) the final answer after converting will be 732.4 mmHg